OpenStudy (sissyedgar):

Which ordered pairs in the form (x, y) are solutions to the equation (x + 6)y = 9? Choose all answers that are correct. A. (−5, 9) B. (−3, 3) C. (3, 1) D. (7, 9)

OpenStudy (sissyedgar):

@misty1212 @Shalante

OpenStudy (sissyedgar):

@ganeshie8

OpenStudy (sissyedgar):

@dan815 @Abhisar

OpenStudy (anonymous):

Make y by itself. Know how?

OpenStudy (sissyedgar):

i do but my brain isnt functionin rn...

OpenStudy (sissyedgar):

:P

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Divide x+6 on both sides \[y=\frac{ 9 }{ x+6 }\] There. You can plug it in.

OpenStudy (sissyedgar):

is B one of em?

OpenStudy (sissyedgar):

i gotta hurry, imma have 2 go 2 bed soon! :( I hate Math! -.-

OpenStudy (anonymous):

Yes, it is.

OpenStudy (sissyedgar):

YAY!

OpenStudy (sissyedgar):

is it 1 more or 2 more left

OpenStudy (anonymous):

One more.

OpenStudy (sissyedgar):

C? Maybe?

OpenStudy (sissyedgar):

no?

OpenStudy (sissyedgar):

well, imma go wit C cuz im bein forced 2 go 2 bed! :(

OpenStudy (anonymous):

It would be A

OpenStudy (sissyedgar):

2 late! :( i hit submit

OpenStudy (anonymous):

9=9/(6-5) 9=9/1 9=9

OpenStudy (anonymous):

Lol why didnt you wait?